# SITE: towards Spatial Intelligence Thorough Evaluation

### v1

@@ 5.3 Spatial Intelligence on Downstream Tasks
To understand why spatial intelligence is important, we conduct a toy experiment where we evaluate multiple VLMs of varying sizes on other real-world embodied tasks using the LIBERO-Spatial [42] dataset. Our goal is to analyze the relationship between performance on spatial intelligence benchmarks and a model’s capability to perform well in real-world tasks. Thus, we finetune and evaluate both variants of the LLaVA-OneVision and Qwen2.5-VL model variants under the few-shot setting. Specifically, we only use 40-160 trajectories from each task in the spatial suite to train each model, but we do not notice much difference. We use a constant learning rate of 2e-5 and finetune each model for 30 epochs. In Table 5, we report the mean L2 distance between the final positions of the target object and robot arm effector across all episodes as well as the overall success rate. The negated mean L2 distance and CAA scores on our SITE benchmark across all VLMs have a positive Pearson Correlation Coefficient of 0.902. Notably, we also observe that the Qwen2.5-VL-7B model achieves a success rate of 38% while the others fail completely. These results suggest that pretraining data recipes and scale are both important for improving spatial intelligence in VLMs. We provide a more detailed correlation analysis in section III of the appendix, comparing the impact of different benchmarks on embodied tasks. Importantly, these results indicate that AI agents have to possess a high degree of spatial intelligence to reason and interact effectively in the physical world.
@@ 6 Conclusion
In this work, we introduce SITE, a comprehensive benchmark focused on evaluating VLMs’ ability to perform visuospatial reasoning. We pull tasks from 31 datasets and design two novel types of tasks for view-taking and dynamic scenarios. Evaluation on SITE demonstrates a huge gap between humans and state-of-the-art VLMs. Moreover, we empirically demonstrate the positive correlation between the performance of VLMs on SITE and a robot manipulation task.
@@ I Datasets Surveyed
  • Blink [19]: a comprehensive benchmark designed to evaluate multimodal large language models(MLLMs) across broad visual perception tasks.
  • CLEVR [30]: a visual question answering dataset contained various aspects of visual reasoning tasks.
  • CVBench [64]: a vision-centric benchmark evaluating 2D and 3D understanding of the models.
  • GQA [28]: a visual question answering benchmark designed to test compositional reasoning and spatial understanding constructed from structured scene representations.
  • IconQA [46]: a dataset targeting diagram-based question answering, challenging models to interpret and reason over abstract visual representations.
  • LogicVista [72]: a benchmark aimed at assessing the logical reasoning capabilities of MLLMs through structured visual tasks.
  • MMBench [44]: a large-scale benchmark for evaluating the performance of multimodal models across a wide range of vision-language tasks.
  • MME [17]: a comprehensive evaluation benchmark for MLLMs, covering various aspects of perception and cognition abilities.
  • MME-RealWorld [83]: a real-world multimodal evaluation benchmark that tests models on practical perception tasks.
  • MMIU [52]: a comprehensive benchmark designed to evaluate multi-images tasks on MLLMs.
  • MMTBench [76]: an evaluation benchmark contained massive multimodal tasks from various scenarios.
  • MMVet [79]: an evaluation suite focusing on the integration of multiple Vision-Language capabilities.
  • MMVP [65]: an MLLM evaluation benchmark focusing on CLIP-blind pairs.
  • MuirBench [68]: a comprehensive benchmark targeting at robust multi-images understanding abilities of multimodal models.
  • SAT [58]: a spatial training dataset with static and dynamic spatial reasoning tasks.
  • SeedBench [36, 35, 37]: an evaluation benchmark for generative comprehension capabilities of MLLMs.
  • SpatialEval [69]: a novel benchmark that covers different aspects of spatial reasoning in textual and visual formats.
  • SPEC [56]: a synthetic dataset designed to test the fine-grained vision-language understanding of models.
  • VQAv2 [34]: an enhanced version of the Visual Question Answering dataset, providing more balanced question-answer pairs to reduce language biases and better evaluate visual understanding.
  • VSR [43]: a benchmark designed to assess visual spatial reasoning capabilities within images.
  • VStarBench [70]: a visual question answering benchmark that focuses on detailed visual grounding on high-resolution images.
  • 3DSRBench [47]: a comprehensive 3D spatial reasoning benchmark on diverse entities.
@@
  • ActivityNetQA [81]: a large-scale video question answering dataset based on ActivityNet, designed to evaluate models’ abilities to comprehend and reason about complex human activities in videos.
  • MLVU [84]: a comprehensive benchmark designed for long video undertsanding
  • MVBench [40]: a comprehensive benchmark for multimodal video understanding, assessing models on a variety of tasks.
  • Open-EQA [48]: an open-ended embodied question answering dataset that tests models’ abilities to interact with and reason about 3D environments through videos and natural language queries.
  • TGIFQA [29]: a dataset for spatiotemporal reasoning in video question answering through tasks like action recognition and repetition counting.
  • TVQA [34]: a video question answering dataset constructed from TV shows, focusing on temporal and contextual reasoning.
  • VideoMME [18]: a comprehensive benchmark for evaluating multimodal large language models on video understanding tasks.
  • VSI-Bench [74]: a benchmark designed to evaluate spatial reasoning abilities of MLLMs, focusing on tasks that require understanding spatial relationships within indoor scenes.
Ego-Exo4D [22]: a large-scale, multimodal, multiview video dataset capturing skilled human activities from synchronized egocentric and exocentric perspectives. It encompasses over 1,286 hours of video data collected from 740 participants across 13 cities, featuring diverse tasks such as cooking, sports, and music. The dataset includes rich annotations like expert commentary, narrate-and-act descriptions, and atomic action labels, supporting benchmarks in fine-grained activity recognition, proficiency estimation, cross-view translation, and pose estimation.
@@ III Dataset Statistics
Besides, we include more dataset statistics, a radar chat visualizing the comparison of various models, more illustrative examples in our dataset, an example about the frame-reordering task.
@@
57(Ifyouareunsurewhichexactcategoryitbelongsto,choose"Other spatial category can not be sure.")
@@
LLaVA-OV-0.5B 0.268 $\pm$ 0.241 0.0 35.9 87.8 67.5 58.3 51.8 62.5 18.4
LLaVA-OV-7B 0.142 $\pm$ 0.172 0.0 62.6 88.4 95.4 62.2 69.9 78.9 30.2
Qwen2.5-VL-3B 0.139 $\pm$ 0.153 0.0 61.2 85.9 81.4 82.8 65.5 74.9 29.5
Qwen2.5-VL-7B 0.030 $\pm$ 0.040 38.0 68.1 85.9 89.0 88.8 68.4 77.7 31.4
Correlation - - 0.935 -0.602 0.749 0.832 0.842 0.847 0.902

### v2

@@ 5.3 Spatial Intelligence on Downstream Tasks
To understand why spatial intelligence is important, we conduct a toy experiment where we evaluate multiple VLMs of varying sizes on other real-world embodied tasks using the LIBERO-Spatial [liu2023libero] dataset. Our goal is to analyze the relationship between performance on spatial intelligence benchmarks and a model’s capability to perform well in real-world tasks. Thus, we fine-tune and evaluate both variants of the LLaVA-OneVision and Qwen2.5-VL model variants under the few-shot setting. Specifically, we only use 40-160 trajectories from each task in the spatial suite to train each model, but we do not notice much difference. We use a constant learning rate of 2e-5 and fine-tune each model for 30 epochs. In Table 5, we report the mean L2 distance between the final positions of the target object and robot arm effector across all episodes as well as the overall success rate. The negated mean L2 distance and CAA scores on our SITE benchmark across all VLMs have a positive Pearson Correlation Coefficient of 0.902. Notably, we also observe that the Qwen2.5-VL-7B model achieves a success rate of 38% while the others fail completely. These results suggest that pretraining data recipes and scale are both important for improving spatial intelligence in VLMs. We provide a more detailed correlation analysis in the appendix, comparing the impact of different benchmarks on embodied tasks. Importantly, these results indicate that AI agents have to possess a high degree of spatial intelligence to reason and interact effectively in the physical world.
@@ 6 Conclusion
In this work, we introduce SITE, a comprehensive benchmark focused on evaluating VLMs’ ability to perform visuospatial reasoning. We pull tasks from 30 existing datasets and then design two novel types of tasks for view-taking and dynamic scenarios. Evaluation on SITE demonstrates a huge gap between humans and state-of-the-art VLMs. Moreover, we empirically demonstrate the positive correlation between the performance of VLMs on SITE and a robot manipulation task.
[…]
@@ I Datasets Surveyed
  • Blink [fu2024blink]: a comprehensive benchmark designed to evaluate multimodal large language models(MLLMs) across broad visual perception tasks.
  • CLEVR [johnson2016clevrdiagnosticdatasetcompositional]: a visual question answering dataset containing various aspects of visual reasoning tasks.
  • CVBench [tong2024cambrian1]: a vision-centric benchmark evaluating 2D and 3D understanding of the models.
  • GQA [hudson2018gqa]: a visual question answering benchmark designed to test compositional reasoning and spatial understanding constructed from structured scene representations.
  • IconQA [lu2022iconqanewbenchmarkabstract]: a dataset targeting diagram-based question answering, challenging models to interpret and reason over abstract visual representations.
  • LogicVista [xiao2024logicvistamultimodalllmlogical]: a benchmark aimed at assessing the logical reasoning capabilities of MLLMs through structured visual tasks.
  • MMBench [liu2024mmbenchmultimodalmodelallaround]: a large-scale benchmark for evaluating the performance of multimodal models across a wide range of vision-language tasks.
  • MME [fu2024mmecomprehensiveevaluationbenchmark]: a comprehensive evaluation benchmark for MLLMs, covering various aspects of perception and cognition abilities.
  • MME-RealWorld [zhang2024mme]: a real-world multimodal evaluation benchmark that tests models on practical perception tasks.
  • MMIU [meng2024mmiu]: a comprehensive benchmark designed to evaluate multi-image tasks on MLLMs.
  • MMTBench [mmtbench]: an evaluation benchmark containing massive multimodal tasks from various scenarios.
  • MMVet [yu2024mm]: an evaluation suite focusing on the integration of multiple Vision-Language capabilities.
  • MMVP [tong2024eyes]: an MLLM evaluation benchmark focusing on CLIP-blind pairs.
  • MuirBench [wang2024muirbenchcomprehensivebenchmarkrobust]: a comprehensive benchmark targeting robust multi-image understanding abilities of multimodal models.
  • SAT [ray2024satspatialaptitudetraining]: a spatial training dataset with static and dynamic spatial reasoning tasks.
  • SeedBench [li2023seed, li2023seed2, li2024seed2plus]: an evaluation benchmark for generative comprehension capabilities of MLLMs.
  • SpatialEval [wang2025picture]: a novel benchmark that covers different aspects of spatial reasoning in textual and visual formats.
  • SPEC [peng2024synthesize]: a synthetic dataset designed to test the fine-grained vision-language understanding of models.
  • VQAv2 [lei2018tvqa]: an enhanced version of the Visual Question Answering dataset, providing more balanced question-answer pairs to reduce language biases and better evaluate visual understanding.
  • VSR [Liu2022VisualSR]: a benchmark designed to assess visual spatial reasoning capabilities within images.
  • VStarBench [vstar]: a visual question answering benchmark that focuses on detailed visual grounding on high-resolution images.
  • 3DSRBench [ma20243dsrbenchcomprehensive3dspatial]: a comprehensive 3D spatial reasoning benchmark on diverse entities.
@@
  • ActivityNetQA [yu2019activityqa]: a large-scale video question answering dataset based on ActivityNet, designed to evaluate models’ abilities to comprehend and reason about complex human activities in videos.
  • MLVU [MLVU]: a comprehensive benchmark designed for long video understanding
  • MVBench [li2024mvbenchcomprehensivemultimodalvideo]: a comprehensive benchmark for multimodal video understanding, assessing models on a variety of tasks.
  • Open-EQA [OpenEQA2023]: an open-ended embodied question answering dataset that tests models’ abilities to interact with and reason about 3D environments through videos and natural language queries.
  • TGIFQA [jang2017tgifqaspatiotemporalreasoningvisual]: a dataset for spatiotemporal reasoning in video question answering through tasks like action recognition and repetition counting.
  • TVQA [lei2018tvqa]: a video question answering dataset constructed from TV shows, focusing on temporal and contextual reasoning.
  • VideoMME [fu2024video]: a comprehensive benchmark for evaluating multimodal large language models on video understanding tasks.
  • VSI-Bench [yang2024thinkingspacemultimodallarge]: a benchmark designed to evaluate spatial reasoning abilities of MLLMs, focusing on tasks that require understanding spatial relationships within indoor scenes.
Ego-Exo4D [grauman2024egoexo4dunderstandingskilledhuman]: a large-scale, multimodal, multiview video dataset capturing skilled human activities from synchronized egocentric and exocentric perspectives. It encompasses over 1,286 hours of video data collected from 740 participants across 13 cities, featuring diverse tasks such as cooking, sports, and music. The dataset includes rich annotations like expert commentary, narrate-and-act descriptions, and atomic action labels, supporting benchmarks in fine-grained activity recognition, proficiency estimation, cross-view translation, and pose estimation.
@@ III Dataset Statistics
Besides, we include more dataset statistics, a radar chart comparing various models, more examples in our dataset, an example about the frame-reordering task, and a bar chart to reveal data decomposition.
@@
57(Ifyouareunsurewhichexactcategoryitbelongsto,choose"Other spatial categories can not be sure.")
@@
Figure 12: Dataset composition for each spatial category.